# Adaptive Convolution for CNN-based Speech Enhancement Models

### v1

@@ III-B Joint Attention
Inspired by ODConv [23], the joint attention mechanism can be further extended to generate attention maps for other dimensions of the kernel space, including the input channel, output channel, and spatial dimensions. They are generated by adding attention heads, followed by sigmoid activation. These optional structures enable more precise adjustment of convolution kernels, enhancing their adaptability to the time-varying speech statistical characteristics.
@@ III-C Multi-frame Parallelism
Adaptive convolution generates independent convolution kernels for each frame in a causal manner, preventing parallel computation across different frames of an utterence. While this limitation does not affect streaming inference, where frames are processed sequentially, it poses a challenge for parallel computing across all frames in an audio sample. In addition to explicitly implementing convolution through matrix multiplication, we propose two approaches to enable multi-frame parallelism.
@@
$Z\left(t\right)=W\left(t\right)\circledast Y\left(t\right)=\left[\sum_{k=1}^{K}A_{k}\left(t\right)W_{k}\right]\circledast Y\left(t\right)$ (5)
$=\sum_{k=1}^{K}A_{k}\left(t\right)\bigg{[}W_{k}\circledast Y\left(t\right)\bigg{]}=\sum_{k=1}^{K}A_{k}\left(t\right)Z_{k}\left(t\right),$
@@ IV AdaptCRN
Adaptive convolution achieves significant performance improvements with negligible increases in computational complexity, making it an effective component for ultra-lightweight models designed for real-time applications, which typically impose strict constraints on computational resources. Additionally, due to the inherently small parameter budgets of ultra-lightweight models, the total number of parameters remains within an acceptable range even when adaptive convolution is incorporated. Therefore, we propose the adaptive convolution recurrent network (AdaptCRN), an ultra-lightweight model that leverages adaptive convolution to enhance performance while maintaining efficiency.
@@ IV-B Spectral Compression
where $X,X_{\text{r}},X_{\text{i}}$ denote the noisy spectrum and its real and imaginary parts, respectively, the superscript c represents the compressed features, and $f_{\text{ERB}}(\cdot)$ denotes the frequency dimension downsampling operation based on the ERB scale. For the compressed features, we employ the SFE module from [7] to integrate information from adjacent subbands into the channel dimension. The spectral decompression module upsamples the frequency dimension of the single-channel output of the decoder to produce the predicted mask. The decompression is achieved by applying the transpose of the frequency dimension downsampling matrix.
@@ IV-C Adaptive Block
Inverted residual blocks, ConvNeXt blocks, and StarNet blocks are modern CNN designs that have been proven to be lightweight and efficient. These blocks share a similar structure comprising one depthwise convolution (DW Conv) layer and two pointwise convolution (PW Conv) layers. Inspired by these designs, we propose the basic block for AdaptCRN, as depicted in Fig. 2(b). The input features are first processed by layer normalization (LN) along the channel and frequency dimensions, which smooths the energy distribution and preserves spectral patterns across different channels and subbands. The element-wise affine transformation in LN acts as a set of fixed filters to preprocess spectral features. Subsequently, the features pass through a DW Conv layer followed by two PW Conv layers in sequence. Both the DW Conv and the second PW Conv are followed by batch normalization (BN) and parametric ReLU (PReLU) activation, while Gaussian error linear unit (GELU) activation is applied between the two PW Conv layers. We also consider the star operation from StarNet as a replacement for GELU, which is defined as $f_{\text{ReLU6}}\left(x\right)*x$, with $x$ the input features and $f_{\text{ReLU6}}(\cdot)$ the ReLU6 activation function. However, experimental results do not indicate a performance advantage over GELU. Additionally, when frequency dimension downsampling is not required, skip connections are employed to directly add the input features to the output features, mitigating information loss in deeper layers.
@@
By incorporating adaptive convolution into the basic block, we derive the adaptive block, as illustrated in Fig 2(c). This block employs joint attention to generate both the input/output temporal channel attention maps and the adaptive kernel attention weights for the three sub-layers. Notably, adaptive convolution introduces high nonlinearity to the model, which is particularly evident when the activation function between the two PW Conv layers is removed. Typically, two PW Conv layers without activation can be reparameterized into a single layer, expressed as
$W^{(2)}\left(W^{(1)}x\right)=\left(W^{(2)}W^{(1)}\right)x,$ (8)
@@
where $W^{(1)},W^{(2)}$ denote the kernels of the two PW Conv layers, respectively, and the bias term is omitted for simplicity, as its inclusion does not alter the result. However, with adaptive convolution, the reparameterization changes to
$\left(\sum_{k=1}^{K}A_{k}^{(2)}W_{k}^{(2)}\right)\left[\left(\sum_{k=1}^{K}A_{k}^{(1)}W_{k}^{(1)}\right)x\right]$ (9)
$=\left[\sum_{j=1}^{K}\sum_{i=1}^{K}A_{j}^{(2)}A_{i}^{(1)}\left(W_{j}^{(2)}W_{i}^{(1)}\right)\right]x,$
@@
making two adaptive PW Conv layers with $K$ candidate kernels equivalent to a single layer with $K^{2}$ kernels. This observation inspires a trade-off between model size and computational complexity. By reducing the number of layers while increasing the number of candidate kernels, we can decrease computational overhead at the expense of increased parameters. Additionally, since the resulting $K^{2}$ kernels and their corresponding kernel attention weights are derived from two groups of $K$ kernels, they reside in a low-rank space relative to directly optimizing $K^{2}$ independent kernels. Consequently, the reparameterized results may exhibit slightly reduced performance compared to directly optimizing a single layer with $K^{2}$ kernels.
@@ IV-D Loss Function
[…]
$\mathcal{L}=\lambda_{1}\mathcal{L}_{\text{SI-SNR}}(\hat{s},s)+\lambda_{2}\mathcal{L}_{\text{mag}}(\hat{S},S)$ (13)
$+\lambda_{3}\left(\mathcal{L}_{\text{real}}(\hat{S},S)+\mathcal{L}_{\text{imag}}(\hat{S},S)\right),$
@@ V-B Baseline Models and Ablation Study
Ablation experiments are conducted on DPCRN-light to varify the effectiveness of the proposed channel modeling techniques and the joint channel and spatial attention mechanism. Additionally, we validate the advantage of frame-level kernel adjustment by comparing with utterance-level global dynamic convolution, which employs non-causal attention with global pooling. We further investigate the impact of expanding the kernel space by increasing the number of candidate kernels, as well as the performance of directly optimizing the kernels.
@@ V-C Implementation Details
The STFT is computed with a segment length of 32 ms, an overlap of 50%, a fast Fourier transform (FFT) length of 512, and a square-root Hanning window for analysis and synthesis. For adaptive convolution, the number of candidate kernels $K$ is set to 8, except in experiments specifically investigating its effect. The number of hidden channels for channel modeling is set to 32, while the convolution kernel size of the 1D-Conv used in multi-frame channel modeling is set to 3. Optional joint channel and spatial attention mechanisms are not used unless specified in ablation experiments.
@@
For AdaptCRN, as detailed in Section IV-B, the amplitude, real part, and imaginary part of the noisy spectrum are fed into the spectral compression module. The first 65 low-frequency bands remain unaltered, while the 192 high-frequency bands are mapped to 64 bands. 9-channel, 129-dimensional features per frame are obtained using SFE with a kernel size of 3. The hidden channels of all adaptive blocks are set to 16, except in the final block, where the hidden channels between the two PW Convs are reduced to 4. In the encoder, the convolution kernel sizes of the first two layers are (1,5) with a frequency dimension stride of 2, while the last three layers use kernel sizes of (3,3) with a stride of 1. In the decoder, the last two layers employ depthwise transposed convolutions, with the remaining settings matching the encoder. The grouped DPRNN consists of 2 groups, with a frequency dimension of 33. The hidden channels of the intra-frame and inter-frame RNNs are set to 8 and 16, respectively. The single-channel output of the final adaptive block undergoes spectral decompression followed by a learnable sigmoid function to generate the amplitude mask.
@@ VI-A1 Ablation study results for kernel attention
Table I presents the results of ablation experiments conducted on DPCRN-light, exploring the configuration of kernel attention for adaptive convolution. In these experiments, the number of candidate convolution kernels is fixed at 8. The table reports trainable parameter numbers, MACs, and evaluation results on DNS5 test set for three channel modeling approaches: single-frame, multi-frame, and temporal channel modeling. The results demonstrate that the adaptive convolution with any channel modeling approach significantly outperforms the vanilla convolution across all metrics, achieving approximately a 0.1 gain in PESQ and a 0.05 improvement in DNSMOS-OVRL, with only a negligible increase in computational complexity. Among the three approaches, temporal channel modeling introduces more parameters but achieves the most substantial performance gains. This indicates that utilizing a GRU for temporal channel modeling effectively leverages historical information, enabling the generation of more suitable adaptive kernels for each frame.
@@
VI-A2 Number of candidate kernels
Table II reports the performance of adaptive convolution with varying numbers of candidate kernels applied to DPCRN-light on DNS5 test set, with $K$ configured as 4, 8, 16, 32, and 64. The results indicate that expanding the kernel space by increasing the number of candidate kernels is an effective strategy to improve performance. This observation provides a flexible approach to trading model parameters for enhanced performance, with only a minimal increase in computational complexity. However, the performance gains diminish as $K$ further increases. Specifically, increasing $K$ from 32 to 64 yields a slight decrease in DNSMOS, although PESQ improves by 0.02. This suggests that the efficiency of parameter utilization decreases as the kernel space expands, and the performance improvement from adaptive convolution may approach saturation.
Additionally, we investigate the scenario of directly optimizing the convolution kernel, where the kernel attention module predicts the full parameters of the convolution kernel rather than the weights for candidate kernels. While this approach may provide greater flexibility, it significantly increases computational complexity due to the large output dimension of the attention module, and its performance remains suboptimal. These results highlight the challenges of directly optimizing convolutional kernels and underscore the advantages of leveraging a weighted sum of candidate kernels.
@@ VI-A2 Number of candidate kernels
Globle dynamic convolution 358.62 194.58 15.060 0.758 2.410 2.924 3.210 4.023
Single-frame channel modeling 358.62 195.44 14.924 0.761 2.390 2.946 3.229 4.034
Multi-frame channel modeling 378.21 196.67 14.956 0.762 2.406 2.957 3.239 4.038
Temporal channel modeling 410.21 198.79 15.147 0.765 2.427 2.964 3.246 4.040
+ Joint channel attention 426.71 199.83 15.177 0.766 2.445 2.974 3.257 4.043
+ Joint spatial attention 412.45 198.93 15.133 0.766 2.427 2.971 3.252 4.045
+ Joint channel/spatial attention 428.95 199.97 15.192 0.766 2.445 2.970 3.253 4.040
TABLE II: Experimental results for the number of candidate convolution kernels on DNS5 test set.
  Number of candidate kernels Para. (K) MACs (M) SI-SNR ESTOI PESQ DNSMOS
@@
$K=4$ 257.59 198.70 14.900 0.759 2.401 2.946 3.227 4.039
$K=8$ 410.21 198.79 15.147 0.765 2.427 2.964 3.246 4.040
$K=16$ 715.44 198.96 15.172 0.766 2.437 2.976 3.258 4.045
$K=32$ 1325.90 199.30 15.278 0.769 2.453 2.982 3.266 4.043
$K=64$ 2546.83 199.99 15.280 0.770 2.472 2.979 3.262 4.042
Directly optimizing the kernels 1353.17 277.25 15.139 0.767 2.451 2.969 3.252 4.042

### v2

@@ III-B Joint Attention
Inspired by ODConv [23], the joint attention mechanism can be further extended to generate attention maps for other dimensions of the kernel space, including joint channel attention and joint spatial attention. Specifically, joint channel attention assigns time-varying attention weights $\mathbf{A}^{\text{c}}(t)\in\mathbb{R}^{C_{\text{in}}}$ and $\mathbf{A}^{\text{f}}(t)\in\mathbb{R}^{C_{\text{out}}}$ to the input channel and output channel dimensions of the kernel. Joint spatial attention generates time-varying weights $\mathbf{A}^{\text{s}}\in\mathbb{R}^{K_{\text{t}}\times K_{\text{f}}}$ for each spatial location of the convolution kernel with shape $(K_{\text{t}},K_{\text{f}})$. For efficiency, all attention weights are shared across the candidate kernels. The kernel aggregation incorporating joint channel and spatial attention is formulated as
$\mathbf{W}\left(t\right)=\mathbf{A}^{\text{s}}\left(t\right)*\mathbf{A}^{\text{c}}\left(t\right)*\mathbf{A}^{\text{f}}\left(t\right)*\left[\sum_{k=1}^{K}A_{k}\left(t\right)\mathbf{W}_{k}\right],$ (5)
where the attention weights are implicitly broadcasted to align with the shape of $\mathbf{W}_{k}$ before element-wise multiplication. These attention weights are generated by extending the kernel attention module with additional attention heads followed by sigmoid activation, similar to the design of joint multi-layer attention. These optional structures enable more precise adjustment of convolution kernels, enhancing their adaptability to the time-varying speech statistical characteristics.
@@ III-C Multi-frame Parallelism
Adaptive convolution generates independent convolution kernels for each frame in a causal manner, preventing parallel computation across different frames of an utterance. While this limitation does not affect streaming inference, where frames are processed sequentially, it poses a challenge for parallel computing across all frames in an audio sample. In addition to explicitly implementing convolution through matrix multiplication, we propose two approaches to enable multi-frame parallelism.
@@
$\mathbf{Z}\left(t\right)=\mathbf{W}\left(t\right)\circledast\mathbf{Y}\left(t\right)=\left[\sum_{k=1}^{K}A_{k}\left(t\right)\mathbf{W}_{k}\right]\circledast\mathbf{Y}\left(t\right)$ (6)
$=\sum_{k=1}^{K}A_{k}\left(t\right)\bigg[\mathbf{W}_{k}\circledast\mathbf{Y}\left(t\right)\bigg]=\sum_{k=1}^{K}A_{k}\left(t\right)\mathbf{Z}_{k}\left(t\right),$
@@ IV AdaptCRN
Adaptive convolution achieves significant performance improvements with limited increases in computational complexity, making it an effective component for ultra-lightweight models designed for real-time applications, which typically impose strict constraints on computational resources. Additionally, in ultra-lightweight models, the inherently small convolution kernel size ensures that both the total parameter count and the added computational load remain within an acceptable range when adaptive convolution is incorporated. Therefore, we propose the adaptive convolution recurrent network (AdaptCRN), an ultra-lightweight model that leverages adaptive convolution to enhance performance while maintaining efficiency.
@@ IV-B Spectral Compression
where $X,X_{\text{r}},X_{\text{i}}$ denote the noisy spectrum and its real and imaginary parts, respectively, the superscript c represents the compressed features, and $f_{\text{ERB}}(\cdot)$ denotes the frequency dimension downsampling operation based on the ERB scale. For the compressed features, we employ the SFE module from [7] to integrate information from adjacent subbands into the channel dimension. The spectral decompression module upsamples the frequency dimension of the single-channel output of the decoder to produce the predicted mask. The decompression is achieved by applying the transpose of the frequency dimension downsampling matrix, rather than using a learnable transformation.
@@ IV-C Adaptive Block
Inverted residual blocks, ConvNeXt blocks, and StarNet blocks are modern CNN designs that have been proven to be lightweight and efficient. These blocks share a similar structure comprising one depthwise convolution (DW Conv) layer and two pointwise convolution (PW Conv) layers. Inspired by these designs, we propose the basic block for AdaptCRN, as depicted in Fig. 2(b). The input features are first processed by layer normalization (LN) along the channel and frequency dimensions, which smooths the energy distribution and preserves spectral patterns across different channels and subbands. The element-wise affine transformation in LN acts as a set of fixed filters to preprocess spectral features. Subsequently, the features pass through a DW Conv layer followed by two PW Conv layers in sequence. Both the DW Conv and the second PW Conv are followed by batch normalization (BN) and parametric ReLU (PReLU) activation, while Gaussian error linear unit (GELU) activation is applied between the two PW Conv layers. We also consider the star operation from StarNet as a replacement for GELU, which is defined as $f_{\text{ReLU6}}\left(x\right)*x$, with $x$ the input features and $f_{\text{ReLU6}}(\cdot)$ the ReLU6 activation function. However, experimental results do not indicate a performance advantage over GELU. Additionally, when frequency dimension downsampling is not required and the number of input and output channels is equal, skip connections are employed to directly add the input features to the output features, mitigating information loss in deeper layers.
@@
By incorporating adaptive convolution into the basic block, we derive the adaptive block, as illustrated in Fig 2(c). This block employs joint attention to generate both the input/output temporal channel attention maps and the adaptive kernel attention weights for the three sub-layers. For efficiency, the joint channel attention maps are applied to the input and output features of the three sub-layers, rather than to each individual kernel within those layers as described in Section III-B.
Adaptive convolution introduces high nonlinearity to the model, which is particularly evident when the activation function between the two PW Conv layers is removed. Typically, two PW Conv layers without activation can be reparameterized into a single layer, expressed as
$\mathbf{W}^{(2)}\left(\mathbf{W}^{(1)}\mathbf{x}\right)=\left(\mathbf{W}^{(2)}\mathbf{W}^{(1)}\right)\mathbf{x},$ (9)
@@
where $\mathbf{W}^{(1)},\mathbf{W}^{(2)}$ denote the kernels of the two PW Conv layers, respectively, and the bias term is omitted for simplicity, as its inclusion does not alter the result. However, with adaptive convolution, the reparameterization changes to
$\left(\sum_{k=1}^{K}A_{k}^{(2)}\mathbf{W}_{k}^{(2)}\right)\left[\left(\sum_{k=1}^{K}A_{k}^{(1)}\mathbf{W}_{k}^{(1)}\right)\mathbf{x}\right]$ (10)
$=\left[\sum_{j=1}^{K}\sum_{i=1}^{K}A_{j}^{(2)}A_{i}^{(1)}\left(\mathbf{W}_{j}^{(2)}\mathbf{W}_{i}^{(1)}\right)\right]\mathbf{x},$
@@
making two adaptive PW Conv layers with $K$ candidate kernels equivalent to a single layer with $K^{2}$ kernels. Since the resulting $K^{2}$ kernels and their corresponding kernel attention weights are derived from two groups of $K$ kernels, they reside in a low-rank space relative to directly optimizing $K^{2}$ independent kernels. Consequently, the reparameterized results may exhibit slightly reduced performance compared to directly optimizing a single layer with $K^{2}$ kernels.
@@ IV-D Loss Function
where $s$ and $\hat{s}$ represent clean and enhanced waveforms, $S$ and $\hat{S}$ are their corresponding spectrograms, the subscripts r, i represent the real and imaginary parts of the spectrograms, respectively, $\langle\cdot,\cdot\rangle$ denotes the inner product operator, and $\text{MSE}(\cdot,\cdot)$ represents the mean squared error (MSE). The overall loss function for model training is given by
$\mathcal{L}=\lambda_{1}\mathcal{L}_{\text{SI-SNR}}+\lambda_{2}\mathcal{L}_{\text{mag}}+\lambda_{3}\left(\mathcal{L}_{\text{real}}+\mathcal{L}_{\text{imag}}\right),$ (14)
@@ V-B Baseline Models and Ablation Study
Ablation experiments are conducted on DPCRN-light to verify the effectiveness of the proposed channel modeling techniques and the joint channel and spatial attention mechanism. Additionally, we validate the advantage of frame-level kernel adjustment by comparing with utterance-level global dynamic convolution, which employs non-causal attention with global pooling. Furthermore, we investigate the impact of kernel generation configurations on performance, focusing on the effectiveness of the temperature annealing training strategy and comparing candidate kernel aggregation via softmax-based attention with directly generating kernel parameters through the attention module.
@@ V-C Implementation Details
The STFT is computed with a segment length of 32 ms, an overlap of 50%, a fast Fourier transform (FFT) length of 512, and a square-root Hanning window for analysis and synthesis. For adaptive convolution, the number of candidate kernels $K$ is set to 8, unless otherwise specified. The number of hidden channels for channel modeling is set to 32, while the convolution kernel size of the 1D-Conv used in multi-frame channel modeling is set to 3. Optional joint channel and spatial attention mechanisms are not used unless specified in ablation experiments. During training, we do not apply the temperature annealing strategy for the softmax function described in Section II-A, except in the specific ablation experiment. This is because our experiments show that it does not enhance kernel optimization in adaptive convolution, as detailed in Section VI-A2.
@@
For AdaptCRN, as detailed in Section IV-B, the amplitude, real part, and imaginary part of the noisy spectrum are fed into the spectral compression module. The first 65 low-frequency bands (below 2 kHz) remain unaltered, while the 192 high-frequency bands (above 2 kHz) are mapped to 64 bands using the ERB-based filter described in Section IV-B. 9-channel, 129-dimensional features per frame are obtained using SFE with a kernel size of 3. The hidden channels of all adaptive blocks are set to 16, except in the final block, where the hidden channels between the two PW Convs are reduced to 4. In the encoder, the convolution kernel sizes of the first two layers are (1,5) with a frequency dimension stride of 2, while the last three layers use kernel sizes of (3,3) with a stride of 1. In the decoder, the last two layers employ depthwise transposed convolutions, with the remaining settings matching the encoder. The grouped DPRNN consists of 2 groups, with a frequency dimension of 33. The hidden channels of the intra-frame and inter-frame RNNs are set to 8 and 16, respectively. The single-channel output of the final adaptive block undergoes spectral decompression followed by a learnable sigmoid function to generate the amplitude mask.
@@ VI-A1 Ablation study results for kernel attention
Table I presents the results of ablation experiments conducted on DPCRN-light, exploring the configuration of kernel attention for adaptive convolution. In these experiments, the number of candidate convolution kernels is fixed at 8. The table reports trainable parameter numbers, MACs, and evaluation results on DNS5 test set for three channel modeling approaches: single-frame, multi-frame, and temporal channel modeling. The results demonstrate that the adaptive convolution with any channel modeling approach significantly outperforms the vanilla convolution across all metrics, achieving approximately a 0.1 gain in PESQ and a 0.05 improvement in DNSMOS-OVRL, with only a modest increase in computational complexity. Among the three approaches, temporal channel modeling introduces more parameters but achieves the most substantial performance gains. This indicates that utilizing a GRU for temporal channel modeling effectively leverages historical information, enabling the generation of more suitable adaptive kernels for each frame.
@@
Globle dynamic convolution 358.62 194.88 15.060 0.758 2.410 2.924 3.210 4.023
Single-frame channel modeling 358.62 210.45 14.924 0.761 2.390 2.946 3.229 4.034
Multi-frame channel modeling 378.21 211.68 14.956 0.762 2.406 2.957 3.239 4.038
Temporal channel modeling 410.21 213.80 15.147 0.765 2.427 2.964 3.246 4.040
+ Joint channel attention 426.71 214.84 15.177 0.766 2.445 2.974 3.257 4.043
+ Joint spatial attention 412.45 213.94 15.133 0.766 2.427 2.971 3.252 4.045
+ Joint channel/spatial attention 428.95 214.98 15.192 0.766 2.445 2.970 3.253 4.040
TABLE II: Ablation study results for the detailed configuration of kernel aggregation on DNS5 test set.
  Para. (K) MACs (M) SI-SNR ESTOI PESQ DNSMOS
@@
Adaptive convolution ($K=8$) 410.21 213.80 15.147 0.765 2.427 2.964 3.246 4.040
Adaptive convolution ($K=32$) 1325.90 271.51 15.278 0.769 2.453 2.982 3.266 4.043
w/ temperature annealing 410.21 213.80 15.091 0.765 2.425 2.963 3.243 4.043
w/o softmax 410.22 213.81 15.086 0.764 2.429 2.964 3.244 4.043
w/o softmax & dimensionality reduction 1353.17 277.25 15.139 0.767 2.451 2.969 3.252 4.042
VI-A2 Detailed configuration of kernel generation
Table II first reports the performance of DPCRN-light with adaptive convolution using 32 candidate kernels. While this achieves a slight improvement over our default configuration of $K=8$, expanding the kernel space comes at the cost of significantly more parameters and higher computational complexity. Based on our experiments, 8 candidate kernels provide a reasonable configuration, offering a good trade-off between performance and efficiency. We further examine the temperature annealing strategy for training kernel attention, as recommended in DyConv for CV tasks (see Section II-A). As shown in Table II, applying this strategy to adaptive convolution yields essentially the same results as without it, suggesting that temperature annealing may not be effective for SE models.
Additionally, we evaluate the kernel generation approach recommended in DGCN, where the output of the final FC layer of the attention module is directly used as the convolution kernel parameters. Conceptually, this is closely related to our approach: for each frame, our attention mechanism produces a $K$-dimensional weight vector, which aggregates the $K$ candidate kernels (equivalent to a $K\times D$ matrix, with $D$ denoting the flattened kernel size) into the final kernel. This process is equivalent to using a $K\times D$ FC layer to generate the kernel parameters. However, there are two important differences. First, the softmax normalization forms probability-like weights, each ranging from 0 to 1 and collectively summing to one. DyConv suggests this constraint facilitates kernel learning. Second, the use of kernel attention weights is equivalent to reducing the hidden feature dimension (set to 32) to $K$ (set to 8) before mapping it to the kernel size, thereby introducing a dimensionality reduction step within the attention module.
Table II also presents the results of removing the softmax constraint (replacing it with PReLU). Performance remains essentially unchanged, indicating that DyConv’s insight regarding the benefit of attention score normalization does not hold in SE tasks. We further test directly using the output of the final FC layer as convolution kernel parameters, i.e., removing both the softmax constraint and the dimensionality reduction to $K$. This greatly increases parameter count and computational cost, but the performance improvement is very limited. This demonstrates that employing a low hidden dimension in the FC layer is an effective architectural choice.
Overall, these results provide deeper insight into the kernel generation, showing that certain conclusions drawn in CV do not transfer well to SE. Notably, we retain the kernel aggregation formulation and the softmax constraint on attention weights, as they provide improved interpretability for the kernel generation of adaptive convolution and support our subsequent analysis of the correlation between kernel allocation and speech features.